INAF queue observing April

Observer: E. Sani, M. Faccini, A. Bongiorno
Telescope Operator: S. Anderson, D. Gonzales-Huerta
Support Astronomer: M. Edwards, D. Thompson
Instrument: LBC - MODS - LUCI img

March 31 - April 1

Cirrus above the orizon (W) at beginning of the operations, with low wind. The initial seeing is 0.7 on the GCS.
During the night the average seeing is 1.0 and the sky does remain mostly clear.
After ~15min of the 8th exposure for the program Annibali_23A, we have a shell RIP of the secondary mirror. While the telescope is tracking, the WFS does not work, we thus pause the script. When the TO solves the problem, we are authorized to resume the script. We verify the shell RIP did not affect the scientific frame.

Project_RUN Objects Filter ExpTime Notes
Antoniucci DDT HH34 H 6min COMPLETE
Antoniucci DDT HH34 FeII 8min COMPLETE
Standard Feige56 dual - airmass 1.4
Standard Feige34 dual - red - airmass 1.04
Annibali 23A NGC4449 dualgrating 6h -
Standard HZ44 dual - airmass 1.3

April 1 - April 2

At the beginning of the night we have thin sky conditions with 50% humidity and wind average speed 10 m/s. Seeing is at 0.8/0.9
At the beginning of the scientific operations the average seeing on the GCS is 0.9.

Project_RUN Objects Filter ExpTime Notes
Standard PG0823 dual - airmass 1.08
Fritz_04D GG190 dual 3h20min COMPLETE
Annibali_23A NGC4449 dual 1h30min the last 45min are affected by bad seeing and thick clouds passing. To be repeated.
Covino_10 CXOJ1348+26 dual 2h average seeing within the specs. Moon illumination 60% at 80deg from the target

April 2 - April 3

At the beginning of the twilight we have clear (close to photometric) sky, with 30% humidity and low wind.
Poor seeing, 1.5, at the beginning of the night operations.
The seeing gets better during the first scientific program, with an average value of 1.0 in both channel.
From 9.00UT the seeing is extremely variable, ranging from 1.0 to 1.5.

Project_RUN Objects Filter ExpTime Notes
Standard SA98 U-Special, B-Bessel, g-Sloan, V-Bessel, r-Sloan, I-Bessel - airmass 1.3
Giallongo_27E CLJ1120 g-Sloan, r-Sloan 1h  
Grazian_30A GOODS-North U-spec, r-Sloan 2h  
Brocato_32D Ursa Minor dwarf B-Bessel, V-Bessel - U-special, I-Bessel 15 min - 45 min  
Brocato_32C Draco dwarf B-Bessel, V-Bessel - U-special, I-Bessel 10 min - 32.5 min  

April 3 - April 4

At the beginning of the twilight we have sky with cirrus on the horizon south-west, with 43% humidity and low wind (8-10 m/s).
Good seeing, 1.1, at the beginning of the night operations.
The seeing gets worse during the first scientific program, with an average value of 1.6 in both channels. We thus switch to MODS
Communication problems with M2. The TO needs to calls Doug Miller. After 25 min of efforts by Doug and Runa Briguglio, we are notified that there is no hope to fix the problem in a reasonable time. We thus decide to switch back to LBC.
The seeing gets better after the second switch and stabilizes around 1.0, decreasing till 0.6.
During the Brocato observation there was elongation in the blue channel. Some dofpia reduce the problem for the B images

Project_RUN Objects Filter ExpTime Notes
Giallongo_27E CLJ1120 g-Sloan, r-Sloan 0.5h Bad seeing. To redo
Grazian_30A GOODS-North U-spec, r-Sloan 3h Seeing 1.0 or better
Brocato_DDT Hercules B-Bessel, r-Sloan - U-special, r-Sloan 17+12 min - 8 min The first Br dataset has elongated images in the blue channel; we re-started it within the specs - to be completed

April 4 - April 5

At the beginning of the twilight we have a completely cloudy sky. The dome remains closed due to rain risk.
Forecasts for the night foreseen bad seeing and partly cloudy, we thus require to use MODS. The crew is using the dome-closed time to replace the BCU grating and run tests for M2.
At the beginning of scientific operations we have low wind, 20% humidity with thick stratified clouds.

Project_RUN Objects Filter ExpTime Notes
Annibali_23A NGC4449 dualgrating 45min COMPLETE
Israel_DDT CATS170113 dualgrating 2h COMPLETE

April 5 - April 6

At sunset we have clear sky and wind gusts between 15 m/s and 20 m/s.

Project_RUN Objects Filter ExpTime Notes
Nascimbeni_DDT HAT-P-12 U-Special F972N20 2h43m30s blue 2h55m30s red COMPLETED
Grazian30A GOODS-North U-scpec r-Sloan 30m  

April 6 - April 7

At sunset we have thin clouds on the horizon, low wind and humidity at 37%.
During the first part of the night we have good seeing ~0.7 and thin cirrus, mostly on the horizon.
After 06.20 UT the sky is mostly clear, with thin cirrus remaining only at the SE horizon.
In the second part of the night we have completely clear sky and good seeing <1.0.

Project_RUN Objects Filter ExpTime Notes
Giallongo_27E CLJ1120 g-Sloan r-Sloan 30min within the seeing specs. thin cirrus passing
Standard SA101 U-spec, B-Bessel, g-Sloan, I-, V-Bessel, r-Sloan   airmass 1.2
Grazian_30A GOODS-North U-spec r-Sloan 3h30min within seeing specs
Standard SA104 U-spec, B-Bessel, g-Sloan, I-, V-Bessel, r-Sloan   airmass 1.7
Brocato_32D Ursa Minor Dwarf U-spec I-Bessel; B-Bessel V-Bessel 45min; 15 min COMPLETE

-- MarcoFaccini - 02 Apr 2013
Topic revision: r6 - 10 May 2013, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback